# xGen-VideoSyn-1: High-fidelity Text-to-Video Synthesis with Compressed Representations

### v1

@@ Abstract
We present xGen-VideoSyn-1, a text-to-video (T2V) generation model capable of producing realistic scenes from textual descriptions. Building on recent advancements, such as OpenAI’s Sora, we explore the latent diffusion model (LDM) architecture and introduce a video variational autoencoder (VidVAE). VidVAE compresses video data both spatially and temporally, significantly reducing the length of visual tokens and the computational demands associated with generating long-sequence videos. To further address the computational costs, we propose a divide-and-merge strategy that maintains temporal consistency across video segments. Our Diffusion Transformer (DiT) model incorporates spatial and temporal self-attention layers, enabling robust generalization across different timeframes and aspect ratios. We have devised a data processing pipeline from the very beginning and collected over 13M high-quality video-text pairs. The pipeline includes multiple steps such as clipping, text detection, motion estimation, aesthetics scoring, and dense captioning based on our in-house video-LLM model. Training the VidVAE and DiT models required approximately 40 and 642 H100 days, respectively. Our model supports over 14-second 720p video generation in an end-to-end way and demonstrates competitive performance against state-of-the-art T2V models. Code will be available at https://github.com/SalesforceAIResearch/xgen-videosyn.
@@ 1 Introduction
Text-to-video (T2V) generation models are designed to create videos that depict both realistic scenes from textual descriptions. These models have garnered significant attention from both academia and industry due to current breakthroughs in the field. Recently, Sora [1] demonstrated that it is possible to generate realistic videos of over one minute in length. Despite such impressive advancements, the most capable video generation models remain proprietary and their details undisclosed. In the meantime, a number of open video generation models have emerged, but their capabilities seem to qualitatively under-perform proprietary models. In this work, our primary objective is to design an effective architecture for T2V with competitive performance compared to the state of the art; examine the associated modeling and training technologies; as well as explore the data collection pipeline.
@@
A crucial component of such design is the dimensionality of the latent space determined by the output of the VAE. A latent space with small dimensionality means that the input pixel information needs to be highly compressed, which makes the reconstruction by diffussion more difficult but computationally less expensive. A latent space with large dimensionality makes reconstruction easier but computationally more expensive. In the case of image generation one can choose larger encoding spaces [2] to facilitate reconstruction quality. However, this trade-off is particularly critical for video generation. If we encode each frame independently using an image VAE [10, 11], a 100 frame video of 720p spatial resolution would translate into a latent space of size 100$\times$4$\times$90$\times$160 containing 360, 000 tokens. This makes both training computationally very expensive and inference slow.
@@
To address this issue, we focus on developing a text-to-video (T2V) generation model based on video-specific VAE and DiT technologies. We adopt a video VAE to achieve effective compression of the video pixel information by reducing both spatial and temporal dimensions. That is, instead of encoding each frame independently, we incorporate both temporal and spatial compression. This significantly decreases the token length, improves the computational cost of training and inference, and facilitates the generation of long videos. Additionally, to further reduce computation during long video encoding, we propose a divide-and-merge strategy. This approach splits a long video into multiple segments, which are encoded individually with overlapping frames to maintain good temporal consistency. With the aid of this advanced video VAE, our xGen-VideoSyn-1 model is able to generate videos with over 100 frames at 720p resolution in an end-to-end manner. Figure 1 shows some example videos generated with our model.
@@
In terms of the diffussion stage, we adopt a video diffusion transformer (VDiT) model that is architecturally similar to Latte [11] and Open-Sora [13]. Our VDiT incorporates transformer blocks with both temporal and spatial self-attention layers. We use ROPE [14] and sinusoidal [15] encodings for spatial and temporal position information. This allows for effective generalization across different lengths, aspect ratios, and resolutions. Moreover, our DiT model is trained on a diverse dataset including 240p, 512$\times$512, 480p, 720p, and 1024$\times$1024 resolutions. The video VAE training takes approximately 40 H100 days, while the DiT model requires around 642 H100 days.
@@
Another crucial aspect of T2V models is the data used for training. Generally, these models require high-quality video-text pairs so that the model can better learn the mapping from the text to video modalities. We address this by designing a data processing pipeline that yields a large quantity of high-quality video-text pairs. Our pipeline involves deduplication, OCR, motion and aesthetics analysis, among other processing steps. It also includes a captioning stage, for which we developed a dense video captioning model that produces an average caption length of 84.4 words. Our full pipeline is automatic and scalable. With this pipeline, we have collected a training dataset with over 13M high-quality video-text pairs.
@@
Figure 2: Our xGen-VideoSyn-1 model consists of VDiT and VidVAE that can compress long sequence video into latent.
@@ 4 Training Data Collection and Processing
As shown at Figure 5, the pipeline operates sequentially as follows: 1) Clipping: Splitting long videos into manageable clips; 2) Deduplication: Removing similar and redundant clips; 3) Motion Detection and Re-Clipping: Analyzing motion dynamics across frames to eliminate static video clips and inconsistent frames. 4) OCR Detection: Identifying and removing clips contaminated with text or watermarks. 5) Aesthetics Scoring: Evaluating and scoring the visual quality of clips. 6) Captioning: Adding descriptive captions to the clips.
@@ 4.1 Video Clipping
After loading raw long videos, we use the PySceneDetect [^2] https://github.com/Breakthrough/PySceneDetect tool to divide them into multiple clips with each clip intended to represent a distinct and clean scene. However, we found that some clips still contain redundant or inconsistent scenes, which we address in subsequent steps.
@@ 4.2 Deduplication
The clipping process can occasionally yield clips that are highly similar to one another. To address this, a de-duplication step is essential to filter out redundant clips. We use ffmpeg [^3] https://ffmpeg.org to extract frames and the clip-as-a-service tool [^4] https://github.com/jina-ai/clip-as-service/tree/main to efficiently extract CLIP features and compute similarity scores between clips. In each duplicate pair, we remove the shorter clip based on a similarity score threshold, $\tau$. Through empirical analysis, we have found that a threshold of $\tau$ = 0.9 is effective for identifying duplicates.
@@ 4.6 Dense Captioning
Captioning Model
Our video captioning model is an extended version of xGen-MM [41]. The model architecture is composed of the following four components: (1) the vision encoder (ViT) taking each frame input, (2) the frame-level tokenizer to reduce the number of tokens, (3) the temporal encoder to build video-level token representations, and (4) the LLM generating output text captions based on such video tokens and text prompt tokens.
Specifically, we use a pretrained ViT-H-14 [42] as the vision encoder, designed to take one single image frame at a time. Perceiver-Resampler [43] is then applied to map such visual tokens into $N=128$ visual tokens per frame. The temporal encoder is implemented with Token Turing Machines (TTM) [44], which is a sequential model capable of taking any number of frames to generate a video-level token representation (e.g., $M=128$ tokens regardless the number of frames). Our use of TTM is similar to its usage in Mirasol3B [45], except that our model uses TTM directly to encode a sequence of image tokens while Mirasol3B uses TTM to encode a sequence of low-level video tokens. We use Phi-3 [46] as our multimodal LLM taking such video tokens in addition to the text prompt tokens. For computational efficiency, the model takes uniformly sampled 4 frames per video. As a result, in our model, ViT first maps a video into approximately $4\times 700$ visual tokens, which is then mapped to $4\times 128$ visual tokens using Perceiver-Resampler, and then to $128$ video tokens using TTM.
@@
Analysis of Captioning Results
We randomly sampled 100k captions. Figure 7 (a) shows the caption length distribution and work cloud for these sampled captions. The average caption length is 84.4 words, which is much longer than other video-language datasets as we know. Additionally, about 87% of captions range from 50 to 120 words.
@@ 4.7 Distributed Data Processing Pipeline
To efficiently orchestrate the six data processing and filtering steps described above with minimal manual intervention and optimal resource utilization, we employ a Distributed Data Processing Pipeline. We use RabbitMQ to manage a video processing pipeline. Each stage is deployed with specific resources and processes video IDs through multiple queues. The pipeline starts by adding video IDs to the initial queue, with each subsequent stage filtering and processing clips based on predefined criteria. Further details can be found in the Section A.1 of Appendix.
@@ 6 Conclusion
Appendix A Data Processing
A.1 Distributed Data Processing Pipeline
@@
To achieve the pipeline, we use the RabbitMQ [^7] https://www.rabbitmq.com as a Task Orchestrator (Figure 11). Each process is a deployment with it’s own resource specification, subscribed to a Task Queue. To trigger the pipeline, we start by pushing the video IDs to the initial queue. This is the only manual step required. Once this is done, the clipping process populates the Similarity Score Queue with the video ID. The Similarity Score deployement, subscribed to the corresponding queue, takes up the task, completes it, and pushes only the de-duplicated clip IDs to the Aesthetic Score queue. The Aesthetic score process, after computation, enqueues the OCR detection Queue with only the IDs of only those clips that meet the threshold. In this fashion, the number of clips that are being processed keeps reducing with each step in the pipeline by skipping the computation for clips that do not meet the passing criteria in the previous steps in the pipeline. In addition to the speed gain by skipping computation for failed clips, we also achieve the speed gain due to pipelining. The time taken for each step to process 1000 videos is given in Tab. 5. Equation 5 shows that the pipelined system is 1.5 times faster than the equivalent sequential system. There is scope to further improve this by speeding up the bottleneck process, as the processing time of the pipelined system is dependent on the time taken by the bottleneck process. As shown in Figure 11, our data collection pipeline includes multiple modules such as clipping, captioning and other process.

### v2

@@ Abstract
We present xGen-VideoSyn-1, a text-to-video (T2V) generation model capable of producing realistic scenes from textual descriptions. We extend the latent diffusion model (LDM) architecture by introducing a video variational autoencoder (VidVAE). Our Video VAE compresses video data spatially and temporally, significantly reducing the length of visual tokens and the computational demands associated with generating long-sequence videos. To further address the computational cost, we propose a divide-and-merge strategy that maintains temporal consistency across video segments. Our Diffusion Transformer (DiT) model incorporates spatial and temporal self-attention layers, enabling robust generalization across different time frames and aspect ratios.We also designed a data collection and processing pipeline, which helped us gather over 13 million high-quality video-text pairs. The pipeline includes steps such as clipping, text detection, motion estimation, aesthetics scoring, and dense captioning based on our xGen-MM video-language model. Training the Video VAE and DiT models required approximately 40 and 642 H100 days, respectively. Our model supports over 14-second 720p video generation in an end-to-end way and demonstrates competitive performance against state-of-the-art T2V models.
@@ 1 Introduction
Text-to-video (T2V) generation models are designed to create videos that depict both realistic scenes from textual descriptions. These models are attracting attention both academia and industry due to recent breakthroughs. Recently, Sora [1] demonstrated that it is possible to generate realistic videos of over one minute in length. Despite such impressive advancements, the most capable video generation models remain proprietary and their details undisclosed. While many open video generation models have surfaced recently, their performance lags behind proprietary models. Our goal with this work is to develop a highly effective architecture for text-to-video (T2V) that rivals existing state-of-the-art models. We examine associated modeling and training technologies, as well as explore the data-collection pipeline.
@@
A crucial component of such design is the dimensionality of the latent space determined by the output of the VAE. A latent space with small dimensionality means that the input pixel information needs to be highly compressed, which makes the reconstruction by diffussion more difficult but computationally less expensive. A latent space with large dimensionality makes reconstruction easier but computationally more expensive. In the case of image generation one can choose larger encoding spaces [2] to facilitate reconstruction quality. However, this trade-off is particularly critical for video generation. If we encode each frame independently using an image VAE [10, 11], a 100 frame video of 720p spatial resolution would translate into a latent space of size 100$\times$4$\times$90$\times$160 that contains $360000$ tokens. This makes both training computationally very expensive and inference slow.
@@
To address this issue, we focus on developing a text-to-video (T2V) generation model based on video-specific VAE and DiT technologies. We introduce a video VAE to achieve effective compression of the video pixel information by reducing both spatial and temporal dimensions. That is, instead of encoding each frame independently, we incorporate both temporal and spatial compression. This significantly decreases the token length, improves the computational cost of training and inference, and facilitates the generation of long videos. Additionally, to further reduce computation during long video encoding, we propose a divide-and-merge strategy. This approach splits a long video into multiple segments, which are encoded individually with overlapping frames to maintain good temporal consistency. With the aid of this advanced video VAE, our xGen-VideoSyn-1 model is able to generate videos with over 100 frames at 720p resolution in an end-to-end manner. Figure 1 shows some example videos generated with our model.
@@
In terms of the diffusion stage, we adopt a video diffusion transformer (VDiT) model that is architecturally similar to Latte [11] and Open-Sora [13]. Our VDiT incorporates transformer blocks with both temporal and spatial self-attention layers. We use ROPE [14] and sinusoidal [15] encodings for spatial and temporal position information. This allows for effective generalization across different lengths, aspect ratios, and resolutions. Moreover, our DiT model is trained on a diverse dataset that includes 240p, 512$\times$512, 480p, 720p, and 1024$\times$1024 resolutions. The video VAE training takes approximately 40 H100 days, while the DiT model requires around 642 H100 days.
@@
Training data is crucial for text-to-video models. These models require high-quality video-text pairs to learn how to map text to video modalities. To address this, we designed a data processing pipeline that yields a large quantity of high-quality video-text pairs. Our process includes removing duplicate data, analysis of aesthetics and motion, optical character recognition (OCR), and other processing steps. The process also captions videos. We developed a video captioning model that creates captions with an average of 84.4 words. Our process is automatic and can be scaled as needed. With this process, we have created a training dataset with over 13 million high-quality video-text pairs.
@@
Figure 2: The core of our xGen-VideoSyn-1 model is a Video DiT (VDiT) module and a new Video VAE (VidVAE) module. The key is that our Video VAE module is able to encode and compress long video sequences into a latent representation during training; as well as reconstruct and decode such latent representation into long and realistic video sequences during inference.
@@ 4 Training Data Collection and Processing
The pipeline operates sequentially, as shown in Figure 5. First, it splits long videos into manageable clips. Then, it removes similar and redundant clips. Next, it analyzes aesthetics and motion dynamics across frames to eliminate static video clips and inconsistent frames. After that, it identifies and removes clips contaminated with text or watermarks. Finally, it evaluates and scores the visual quality of clips before adding descriptive captions to the clips. We further describe these steps in this section.
@@ 4.1 Video Clipping
Original long videos are cut into multiple shorter clips using the PySceneDetect [^2] https://github.com/Breakthrough/PySceneDetect tool. Each clip is intended to represent a distinct and clean scene. However, some clips may still contain redundant or inconsistent scenes. These cases are addressed in subsequent steps.
@@ 4.2 Deduplication
The clipping process can sometimes yield clips that are highly similar to one another. To address this, a de-duplication step is essential to filter out redundant clips. We use ffmpeg [^3] https://ffmpeg.org to extract frames and the clip-as-a-service tool [^4] https://github.com/jina-ai/clip-as-service/tree/main to efficiently extract CLIP features and compute similarity scores between clips. In each duplicate pair, we remove the shorter clip based on a similarity score threshold, $\tau$. Through empirical analysis, we have found that a threshold of $\tau$ = 0.9 is effective for identifying duplicates.
@@ 4.6 Dense Captioning
Captioning Model. Our video captioning model is an extended version of xGen-MM [41]. The model architecture is composed of the following four components: (1) a vision encoder (ViT) taking each frame input, (2) a frame-level tokenizer to reduce the number of tokens, (3) a temporal encoder to build video-level token representations, and (4) a LLM generating output text captions based on such video tokens and text prompt tokens.
Specifically, we use a pretrained ViT-H-14 [42] as the vision encoder, designed to take one single image frame at a time. Perceiver-Resampler [43] is then applied to map such visual tokens into $N=128$ visual tokens per frame. The temporal encoder is implemented with Token Turing Machines (TTM) [44], which is a sequential model capable of taking any number of frames to generate a video-level token representation (e.g., $M=128$ tokens regardless the number of frames). Our use of TTM is similar to its usage in Mirasol3B [45], except that our model uses TTM directly to encode a sequence of image tokens while Mirasol3B uses TTM to encode a sequence of low-level video tokens. We use Phi-3 [46] as our multimodal LLM taking such video tokens in addition to the text prompt tokens. For computational efficiency, the model takes uniformly sampled 4 frames per video. Our model uses ViT to map a video into around $4\times 700$ visual tokens. These visual tokens are then mapped to $4\times 128$ visual tokens using Perceiver-Resampler and then to 128 video tokens using TTM.
@@
Analysis of Captioning Results. We randomly sampled 100k captions. Figure 7 (a) shows the caption length distribution and work cloud for these sampled captions. The average caption length is 84.4 words, which is much longer than other video-language datasets as we know. Additionally, about 87% of captions range from 50 to 120 words.
@@ 4.7 Distributed Data Processing Pipeline
To efficiently orchestrate the six data processing and filtering steps described above with minimal manual intervention and optimal resource utilization, we employ a Distributed Data Processing Pipeline. We use RabbitMQ to manage a video processing pipeline. Each stage is deployed with specific resources and processes videos through multiple queues. The pipeline starts by adding videos to the initial queue, with each subsequent stage filtering and processing clips based on predefined criteria. Further details can be found in the Section A of Appendix.
@@ 6 Conclusion
Appendix A Distributed Data Processing Pipeline
@@
To achieve the pipeline, we use RabbitMQ [^7] https://www.rabbitmq.com as a Task Orchestrator (Figure 11). Each process is a deployment with it’s own resource specification, subscribed to a Task Queue. To trigger the pipeline, we start by pushing the video IDs to the initial queue. This is the only manual step required. Once this is done, the clipping process populates the Similarity Score Queue with the video ID. The Similarity Score deployement, subscribed to the corresponding queue, takes up the task, completes it, and pushes only the de-duplicated clip IDs to the Aesthetic Score queue. The Aesthetic score process, after computation, enqueues the OCR detection Queue with only the IDs of only those clips that meet the threshold. In this fashion, the number of clips that are being processed keeps reducing with each step in the pipeline by skipping the computation for clips that do not meet the passing criteria in the previous steps in the pipeline. In addition to the speed gain by skipping computation for failed clips, we also achieve the speed gain due to pipelining. The time taken for each step to process 1000 videos is given in Tab. 5. Equation 5 shows that the pipelined system is 1.5 times faster than the equivalent sequential system. There is scope to further improve this by speeding up the bottleneck process, as the processing time of the pipelined system is dependent on the time taken by the bottleneck process. As shown in Figure 11, our data collection pipeline includes multiple modules such as clipping, captioning and other process.